# Music Foundation Model as Generic Booster for Music Downstream Tasks

### v2

@@ F.3 Results: Mix-Wave-U-Net and CRAFx2 with SoniDo
Incorporating MusicGen features generally leads to improvement over the default network, although not as much as SoniDo. MusicGen Small performed slightly better than the large model, confirming the trend observed in previous downstream tasks. However, improvement over the default network is not always the case with CRAFx2. In the MDXDB21-dry test set, the default model performed better than MusicGen in terms of spectral and panning features as well as stereo-invariant loss. Yet, MusicGen did enhance the dynamic and loudness features. The performance gap of MusicGen in spectral-related metrics, and generally when compared with SoniDo, might stem from MusicGen being trained on $32$-kHz audio, while both our evaluation datasets and training dataset of SoniDo correspond to $44.1$-kHz audio.

### v3

@@ F.3 Results: Mix-Wave-U-Net and CRAFx2 with SoniDo
Incorporating Jukebox and MusicGen features generally leads to improvement over the default network, although not as much as SoniDo. MusicGen Small performed slightly better than the large model, confirming the trend observed in previous downstream tasks. However, for both MusicGen and Jukebox, improvement over the default network is not always the case with CRAFx2. In the MDXDB21-dry test set, the default model performed better than MusicGen in terms of spectral and panning features as well as stereo-invariant loss. Yet, MusicGen did enhance the dynamic and loudness features, while Jukebox only improved loudness and generally led to overfitting. The performance gap of MusicGen in spectral-related metrics, and generally when compared with SoniDo, might stem from MusicGen being trained on $32$-kHz audio, while both our evaluation datasets and training dataset of SoniDo correspond to $44.1$-kHz audio.